OpenStudy (anonymous):

An equation is shown below: 8x + 2(x - 7) = 7x + 3x - 14 Part A: Solve the equation and write the number of solutions. Show all the steps. Part B: Name one property you used to solve this equation.

OpenStudy (anonymous):

Okay, so.... Let's distribute that 2 first... 8x + 2(x - 7) = 7x + 3x - 14 8x + 2x -14 = 7x + 3x -14 10x - 14 = 10x - 14 From here, we see that when we plug in any value of x, it satisfies the equation. Thus all values of x satisfy the given equation. We used the distributive property for this equation.